### v1

# Dataforge: A Data Agent Platform for Autonomous Data Engineering

Abstract
The growing demand for AI applications in fields such as materials discovery, molecular modeling, and climate science has made data preparation an important but labor-intensive step. Raw data from diverse sources must be cleaned, normalized, and transformed to become AI-ready, while effective feature transformation and selection are essential for efficient training and inference. To address the challenges of scalability and expertise dependence, we present Data Agent, a fully autonomous system specialized for tabular data. Leveraging large language model (LLM) reasoning and grounded validation, Data Agent automatically performs data cleaning, hierarchical routing, and feature-level optimization through dual feedback loops. It embodies three core principles: automatic, safe, and non-expert friendly, which ensure end-to-end reliability without human supervision. This demo showcases the first practical realization of an autonomous Data Agent, illustrating how raw data can be transformed “From Data to Better Data.”
I Introduction
Modern AI applications increasingly rely on AI-ready data—data that are clean, consistent, and semantically aligned with model objectives. In fields such as materials discovery, molecular design, and climate modeling, AI-ready data determine whether models can achieve reliable prediction, efficient training, and meaningful scientific insight [yu2025ai]. Yet, transforming raw data into such AI-ready form remains a persistent challenge. Raw datasets are often heterogeneous, incomplete, and noisy, collected from instruments or repositories with inconsistent formats and semantics. Converting them into usable inputs requires extensive cleaning, normalization, and feature transformation, which are still largely manual, error-prone, and dependent on domain expertise. This growing mismatch between the rapid progress of AI models and the slow, fragmented process of data preparation exposes a fundamental data engineering gap that increasingly limits the scalability and reliability of AI applications.
The rapid progress of large language models (LLMs) has opened a new path toward autonomous data engineering. LLMs not only capture extensive domain knowledge from large-scale text corpora but also exhibit emerging abilities to interpret data schemas, infer transformation logic, and generate executable operations—tasks traditionally reserved for human experts. Although many recent works have explored using LLMs for data-centric tasks such as LLM-based feature engineering and transformation [han2024large, hollmann2023large, horn2019autofeat, zhang2023openfe, gong2025evolutionary], they remain limited in several aspects: (1) Lack of full automation: Most approaches still require user intervention, manual code review, or step-by-step prompting to complete a data pipeline. (2) Limited reliability: LLMs may generate invalid or unsafe operations without schema verification or grounding checks, leading to broken workflows. (3) Expert dependency: Many systems assume the user understands LLM prompts, data schemas, or tool APIs, which prevents non-expert adoption.
Newer LLMs now demonstrate emergent agentic behaviors—the ability to reason, plan, and act through iterative interaction with tools and environments. This shift marks an agentic evolution in how intelligence is deployed: systems no longer just generate outputs for a descriptive query, but autonomously understand tasks, reason, plan, ground, execute actions, and adapt through feedback. Extending this paradigm to the data domain gives rise to the concept of Data Agents: autonomous systems that can understand datasets, reason about transformations, and execute data-centric operations to improve quality, consistency, and utility.
Figure 1: Conceptual comparison between a traditional manual workflow and the agentic workflow.
Figure 2: The framework of the Dataforge system.
To this end, we develop the Dataforge, an autonomous data agent specialized for tabular data that embodies three core design goals: 1) Automatic: performs end-to-end data transformation without manual intervention; 2) Safe: ensures workflow reliability via routing, grounding, and execution validation; 3) Non-expert Friendly: allows users with minimal technical background to improve their data by simply specifying task and data types. The dataagent is implemented based on the perception–planning–grounding–execution principles [fu2025autonomous] as a deployable system for tabular data. The data agent includes a sequential stages of data cleaning, hierarchical routing, grounding validation, and dual-loop optimization. In particular, the hierarchical routing mechanism determines both the high-level task family and the low-level feature actions; each action is checked before execution to ensure safety, and results are evaluated to guide iterative improvement; the dual feedback loops ensure that each decision is validated before execution and that performance improvements are optimized through experience replay. Our Contribution: present the first functional data agent demo that leverages LLMs to automatically reason over tabular data and iteratively orchestrate next grounded actions with autostop to fabric AI-ready data.
II Dataforge: Agentic Data Engineering
II-A Overview of the Dataforge
Fig. 2 shows that Dataforge aims to automate the end-to-end data preparation process—from raw input to adaptive refinement—through a structured six-stage pipeline. (1) Input and Cleaning: Raw data is standardized via schema alignment, missing-value handling, and type correction under preset modes (light, aggressive, or time-series). (2) Task-Level Routing: A rule-based router identifies the task type (e.g., classification, clustering, regression) and dispatches data to corresponding action-level LLM router. (3) Action-Level Routing: Within each branch, an LLM-based planner selects, transforms, or generates features based on metadata and experience. (4) Grounding and Execution: Actions are validated for schema and logic safety, then executed through modular Python or SQL engines. (5) Evaluation and Feedback: An evaluator computes task-specific metrics (e.g., F1, AUC, RMSE) and maintains two feedback loops to regulate the process, in which the inner loop validates actions to prevent workflow failure, and the outer loop optimizes performance across iterations. (6) Experience Replay and Summary: The system logs all operations as experiences to replay for continual learning and automatically generate reports for traceable provenance.
II-B Hierarchical Routing
After data cleaning, to enable efficient and reliable decision-making, we adopt a hierarchical routing architecture, including task-level and action-level reasoning. At the task-level routing, a rule-based router quickly identifies the task type: classification, regression, or unsupervised learning, based on table schema metadata, such as, data types, label structures, and feature distribution. Such lightweight router relies on deterministic heuristics, instead of large language models, thus, enable fast and reliable responses across diverse datasets. At the action-level routing, a compact LLM-based planner refines the decision by selects and plans the most suitable feature-level actions such as, different ordered combinations of feature selection, transformation, or generation, under the identified task (e.g., a classification dataset). Since each router operates within a smaller, well-defined action space, this hierarchical routing approach not only accelerates processing but also avoid invalid or high-risk operations.
II-C Dual Feedback Loops
We develop two collaborative feedback loops to transform the static workflow into an adaptive, self-correcting process, in order to achieve autonomy and continual refinement. 1) Action Validation Loop for Safety: This feddback loop is to ground actions to ensure operational safety before execution. Each planned action is first grounded through schema alignment, type checking, and logical consistency tests, such as, detecting divisions by zero or invalid type conversions. Only actions that pass validation proceed to execution so as to prevent runtime errors and maintaining workflow integrity. 2) Performance Optimization Loop for Adaptiveness: This feedback loop focuses on post-execution improvement and enabling the Dataforge to converge toward optimal behavior without human supervision. In particular, after each valid action runs, an evaluator measures task-specific metrics (e.g., F1, AUC, RMSE). If performance improves, the system updates its optimal configuration; otherwise, it triggers re-planning of new data engineering actions. The two loops establish a closed-cycle control mechanism for both grounded safety and self-optimizing adaptivity.
II-D Automatic Summary and Reporting
To ensure transparency and provenance, the system automatically log the routing choices, all actions, grounding outcomes, machine learning performance metrics and summarize these logs into a structured concise natural language report that outlines: (i) the cleaning and feature operations applied, (ii) the evolution of routing decisions, (iii) the validation checks performed, and (iv) the performance progression across iterations. The summary, viewable through the demo interface, translates internal reasoning into a human-readable narrative that lets non-experts trace data transformations and decisions, but also allows the feedback loop to collect experiences as benchmark data and perform self-reflection.
II-E User Interface
Fig. 3 illustrates the user interface of Dataforge, which offers an intuitive environment for autonomous tabular data transformation. Users can upload datasets in standard formats (.csv, .parquet, and .hdf), label predictors and response column in a machine learning task, and configure runtime parameters, such as, Max Actions, Min Improvement, and Time Budget. After uploading datasets, users can select from preloaded examples. During execution, the interface visualizes real-time logs of thinking steps in reasoning, data cleaning and feature engineering actions, and predictive performance metrics (e.g., F1-score, AUC, MAE, and RMSE depending on classification, regression, and unsupervised tasks ), so that users can observe how the data agent progressively refines data quality. Upon completion, it summarizes all executed actions, validation outcomes, and final performance statistics, to provide transparency and provenance, as well as collect action historical trajectories for benchmark data of training data agents. The interface is designed to be modular and extensible, supporting future integration with interactive dashboards and collaborative agent environments.
Figure 3: The interface of Dataforge.
III Demonstration Scenarios
We demonstrate the functionality and effectiveness of Dataforge through both quantitative evaluation and visual inspection of its autonomous workflow. The goal is to show that Dataforge achieves reliable performance across diverse datasets while providing interpretable and user-friendly automation for non-expert users.
III-A Demonstration Setup
All demonstrations are conducted on Ubuntu 22.04.3 LTS with an Intel(R) Core(TM) i9-13900KF CPU and an NVIDIA GeForce RTX 4090 GPU (24 GB, CUDA 12.2). The system is implemented in Python 3.12 using PyTorch for GPU-accelerated data transformation.
III-B Quantitative Evaluations
Study of Overall Performances. TABLE I summarizes the overall performance of Dataforge compared with representative baselines, including classical AutoML methods, RL-based pipelines, and pure LLM approaches. Across all nine datasets, Dataforge achieves the highest or comparable predictive performance while maintaining full automation and zero training cost. Its rule-based task routing and LLM-driven action selection allow it to adapt to different task types efficiently, avoiding unstable or invalid operations common in purely generative systems.
Notably, the agent demonstrates consistent improvements over the baseline data across multiple datasets, validating its ability to autonomously refine data representations through feature transformation and selection. The performance gains are achieved with minimal user input—only specifying the data file, task type, and optional runtime constraints—highlighting the practicality of the proposed autonomous data-processing paradigm.
TABLE I: Downstream performance comparison on regression and classification datasets. For regression, we report 1-RAE (Relative Absolute Error), and for classification, F1-score (%).
Dataset German Credit Amazon Employee Ionosphere PimaIndian Messidor Feature SVMGuide3 OpenML 586 OpenML 618 Airfoil
Source UCIrvine Kaggle UCIrvine Kaggle UCIrvine LibSVM OpenML OpenML UCIrvine
Task Classification Classification Classification Classification Classification Classification Regression Regression Regression
Samples 1,000 32,769 351 768 1,151 1,243 1,000 1,000 1,503
Features 24 9 34 8 19 21 25 50 5
Original 74.20% 93.37% 93.37% 80.68% 69.09% 81.85% 0.6311 0.4402 0.5749
RDG 68.01% 92.31% 91.17% 76.04% 62.38% 78.68% 0.5681 0.3720 0.5193
LDA 63.91% 91.64% 65.53% 63.80% 47.52% 65.24% 0.1109 0.0521 0.2201
ERG 74.43% 92.43% 92.02% 76.17% 66.90% 82.62% 0.6147 0.3561 0.5193
NFS 68.67% 93.21% 91.17% 74.87% 63.77% 79.16% 0.5443 0.3473 0.5193
AFAT 68.32% 92.97% 92.87% 76.56% 66.55% 79.49% 0.5435 0.2472 0.5210
PCA 67.92% 92.29% 92.87% 63.80% 67.21% 67.60% 0.1109 0.1016 0.2730
TTG 64.51% 92.79% 90.31% 74.48% 66.46% 79.81% 0.5443 0.3467 0.5003
FeatLLM [han2024large] 76.35% 93.62% 95.38% 89.66% 72.62% 81.17% 0.6477 0.4597 0.5587
CAAFE [hollmann2023large] 59.92% 91.41% 92.84% 79.86% 66.10% 81.74% N/A N/A N/A
AutoFeat [horn2019autofeat] 74.86% 93.29% 93.37% 80.86% 69.08% 82.54% 0.6329 0.4407 0.5746
OpenFE [zhang2023openfe] 74.50% 93.44% 93.37% 80.86% 69.09% 83.05% 0.6311 0.4402 0.5746
ELLM-FT [gong2025evolutionary] 76.39% 93.17% 96.01% 89.66% 74.80% 82.70% 0.6328 0.4734 0.6174
Pure LLM (Llama 3.1) 75.43% 75.43% 96.68% 89.66% 75.61% 81.17% 0.7811 0.6653 0.6329
RL-Policy-1 (GRFG [wang2022group]) 68.29% 68.29% 93.16% 75.39% 69.24% 79.81% 0.5768 0.4562 0.5587
RL-Policy-2 (MOAT [wang2023reinforcement]) 72.44% 72.44% 95.69% 80.73% 73.02% 81.17% 0.6251 0.4734 0.5967
DataAgent 79.60% 94.41% 97.14% 90.14% 76.98% 84.64% 0.7849 0.7219 0.7492
Study of Reliability and Automation. In practice, Dataforge showed stable behavior. It finished all nine datasets without interruption and usually converged within two action retries on average. Compared with Llama 3.1 8B and Llama 3 8B, which often fail to reach valid results at a 3–5% rate, the agent maintained smooth execution across runs. The average score reached 0.783, slightly higher than both baselines, and no invalid operations were observed. This suggests that the built-in routing and grounding steps make the workflow more predictable and robust for continuous operation.
Study of Efficiency Trade-offs. Compared with reinforcement-learning-based methods, Dataforge runs efficiently without any training cost. RL-Policy-1 and RL-Policy-2 spent over 1,400 s and 650 s on training, respectively, while the agent required none. During inference, the agent finished each workflow in about 3.9 s, with only two calls on average, whereas the RL-based baselines needed 13–22 calls per task. Even with this lightweight process, the agent still achieved the best performance (0.7832). These results show that the system keeps both accuracy and responsiveness, avoiding the heavy overhead of reinforcement-style optimization.
III-C Toward Accurate and Efficient Heart-Disease Detection
Figure 4: Dataforge Dealing with Heart-Disease Detection.
Accurate identification of heart disease is one of the central challenges in clinical cardiology. Modern diagnostic models rely on dozens of imaging and physiological indicators, but these measurements are often noisy, redundant, or inconsistent across patients. As a result, even experienced cardiologists face difficulty in building reliable predictive models from raw data, and existing computational tools usually require manual data engineering and expert parameter tuning.
To address this challenge, DataForge is applied to an automated diagnostic classification task using the SPECTF Heart dataset, which contains patient-level imaging signals used for detecting abnormal cardiac function. The system autonomously analyzes the uploaded dataset, identifies the most informative diagnostic indicators, and removes irrelevant or conflicting variables. It then constructs refined combinations of cardiac features that better represent patient conditions, iteratively evaluating improvements until no further diagnostic gain can be achieved.
Throughout this process (Fig. 4), no manual intervention or algorithmic knowledge is needed. The workflow completes in about 100 seconds and generates a concise report and a new dataset that achieves higher diagnostic accuracy with fewer features, reducing the total number of variables from 44 to 20 while improving the predictive score from 0.772 to 0.840. This demonstrates that DataForge can serve as an intelligent assistant for cardiology, automatically transforming complex measurements into a cleaner, smaller, and more reliable dataset for accurate and efficient heart-disease detection.