### v2

# Dataforge: Agentic Platform for Autonomous Data Engineering

Abstract.
The growing demand for artificial intelligence (AI) applications in materials discovery, molecular modeling, and climate science has made data preparation a critical but labor-intensive bottleneck. Raw data from diverse sources must be cleaned, normalized, and transformed to become AI-ready, where effective feature transformation and selection are essential for robust learning. We present Dataforge, an LLM-powered agentic data engineering platform for tabular data that is automatic, safe, and non-expert friendly. It autonomously performs data cleaning and iteratively optimizes feature operations under a budgeted feedback loop with automatic stopping. Across tabular benchmarks, it achieves the best overall downstream performance; ablations further confirm the roles of routing/iterative refinement and grounding in accuracy and reliability. Dataforge demonstrates a practical path toward autonomous data agents that transform raw data from data to better data.
Data Agent, LLMs, Data-Centric Agentic AI
copyright: none
1. Introduction
Modern AI applications increasingly rely on AI-ready data that are clean, consistent, and semantically aligned with the objectives of the model. In fields such as materials discovery, molecular design, and climate modeling, AI-ready data determine whether models can achieve reliable prediction, efficient training, and meaningful scientific insight (Yu, 2025). However, transforming raw data into such an AI-ready form remains a persistent challenge (Wang et al., 2025b; Ying et al., 2025c). Raw datasets are often heterogeneous, incomplete, and noisy, collected from instruments or repositories with inconsistent formats and semantics. Converting them into usable inputs requires extensive cleaning, normalization, and feature transformation, which are still largely manual, error-prone, and dependent on domain expertise. This growing mismatch between the rapid progress of AI models and the slow and fragmented data preparation process exposes a fundamental data engineering gap that increasingly limits the scalability and reliability of AI applications.
The rapid progress of large language models (LLMs) has opened a new path toward autonomous data engineering. LLMs not only capture extensive domain knowledge from large-scale text corpora but also exhibit emerging abilities to interpret data schemas, infer transformation logic, and generate executable operations—tasks traditionally reserved for human experts (Xie and Ju, 2025). Although many recent works have explored using LLMs for data-centric tasks such as LLM-based feature engineering and transformation (Han et al., 2024; Hollmann et al., 2023; Horn et al., 2019; Zhang et al., 2023; Gong et al., 2025c), they remain limited in several aspects: (1) Lack of full automation: Most approaches still require user intervention, manual code review, or step-by-step prompting to complete a data pipeline. (2) Limited reliability: LLMs may generate invalid or unsafe operations without schema verification or grounding checks, leading to broken workflows. (3) Expert dependency: Many systems assume the user understands LLM prompts, data schemas, or tool APIs, which prevents non-expert adoption.
Newer LLMs now demonstrate emergent agentic behaviors: the ability to reason, plan, and act through iterative interaction with tools and environments. This shift marks an agentic evolution in how intelligence is deployed: systems no longer just generate outputs for a descriptive query, but autonomously understand tasks, reason, plan, ground, execute actions, and adapt through feedback. Extending this paradigm to the data domain gives rise to the concept of Data Agents: autonomous systems that can understand datasets, reason about transformations, and execute data-centric operations to improve quality, consistency, and utility.
Figure 1. Conceptual comparison between a traditional manual workflow and the agentic workflow.
To this end, we develop the Dataforge, an autonomous data agent platform specialized for tabular data that embodies three core design goals: 1) Automatic: performs end-to-end data transformation without manual intervention; 2) Safe: ensures workflow reliability via routing, grounding, and execution validation; 3) Non-expert Friendly: allows users with minimal technical background to improve their data by simply specifying task and data types. The Dataforge is implemented based on the planning–grounding–execution principles (Fu et al., 2025) as a deployable system for tabular data. The data agent includes a sequential set of stages of data cleaning, hierarchical routing, grounding validation, and dual-loop optimization. In particular, the hierarchical routing mechanism determines both the high-level task family and the low-level feature actions; each action is checked before execution to ensure safety, and results are evaluated to guide iterative improvement; the dual feedback loops ensure that each decision is validated before execution and that performance improvements are optimized through experience replay.
Our empirical study demonstrates that Dataforge provides strong and robust gains across diverse tabular datasets: it achieves the best overall downstream utility on six classification and three regression benchmarks compared with classical, RL-based, and LLM-based approaches. Ablation results further verify that hierarchical routing and evaluation-driven refinement are key to performance, while grounding substantially improves executability by reducing invalid actions and repeated retries. It also avoids RL-style training cost and achieves strong performance with low call complexity, supporting its practicality as an automated data engineering tool.
Our Contribution: We present Dataforge, an LLM-powered agentic data engineering platform for tabular data that is automatic, safe, and non-expert friendly, iteratively orchestrating feature operations with automatic stopping to produce AI-ready data.
2. Preliminaries
2.1. Problem Formulation
We consider automated data preparation and feature engineering for tabular datasets. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote a dataset with a feature table $X\in\mathbb{R}^{n\times d}$ (mixed numeric/categorical/time columns) and an optional target column $y$. Given $\mathcal{D}$ and a task specification $\mathcal{T}$ (e.g., classification or regression, evaluation metric, and optional constraints), our goal is to construct a transformation pipeline $\pi$ that maps the raw dataset to an AI-ready representation:
$\pi:\mathcal{D}\rightarrow\mathcal{D}^{\prime}.$
A pipeline $\pi$ is a sequence of executable feature operations $\pi=[a_{1},\ldots,a_{K}]$, where each action $a_{k}$ is selected from a tool library $\mathcal{A}$ (e.g., missing-value imputation, encoding, scaling, feature generation/selection). Applying $\pi$ produces a transformed dataset $\mathcal{D}^{\prime}$ and an execution log. We optimize $\pi$ under a user-defined budget $\mathcal{B}$ (e.g., time limit, maximum number of actions) to improve downstream performance while preserving correctness and safety:
$\max_{\pi}\;\mathrm{Perf}(\mathcal{D}^{\prime},\mathcal{T})\quad\text{s.t.}\quad\mathrm{Safe}(\pi,\mathcal{D})=1,\;\;\mathrm{Cost}(\pi)\leq\mathcal{B}.$
Here, $\mathrm{Perf}(\cdot)$ is measured by a standard metric (e.g., F1 for classification, 1-RAE for regression) using a fixed evaluation protocol, and $\mathrm{Safe}(\cdot)$ indicates whether the pipeline satisfies grounding checks.
2.2. Platform Value
End-to-end automation: Given a dataset and a minimal task specification, Dataforge automatically profiles data, proposes feature operations, validates them, executes transformations, and iteratively improves data utility under a budget. Reliability and safety: Unlike prompt-only approaches, Dataforge enforces action grounding and guardrails to prevent invalid or unsafe transformations and to ensure stable execution. Benchmarkable and reproducible outputs. Dataforge exports a replayable pipeline together with structured logs and summary reports. Non-expert usability: Dataforge exposes task-level controls rather than requiring users to craft prompts or write feature code, lowering the barrier for non-expert practitioners.
2.3. System I/O Contract
Inputs: (i) a tabular dataset (CSV/Parquet or a database table), (ii) an optional target column name, (iii) a task specification $\mathcal{T}$ including task type and evaluation metric, and (iv) optional constraints $\mathcal{B}$ such as time budget, maximum number of actions, and minimum improvement threshold for early stopping.
Outputs: (i) a transformed dataset $\mathcal{D}^{\prime}$ (AI-ready table), (ii) a replayable pipeline $\pi$ that specifies the ordered list of feature operations with parameters, (iii) execution logs that record dataset diffs and validation outcomes for each step, and (iv) a summary report that explains key transformations and the measured utility improvements.
Failure handling: If an action cannot be validated or executed, Dataforge returns a failure reason (e.g., schema mismatch, type violation, suspected leakage, or runtime error) and either (a) replans with an alternative action, or (b) terminates with the best valid pipeline found so far under the budget.
2.4. Important Concepts
AI-ready data: We use AI-ready to describe tabular data that are schema-consistent, type-valid, and suit to model training (e.g., reduced missingness and improved downstream performance).
Operation Set: To refine the feature space, we need to apply mathematical operations to existing features to generate new informative features. All operations are collected in an operation set, denoted by $\mathcal{O}$. These operations can be classified as unary and binary operations. The unary operations such as “square”, “exp”, “log”, etc. The binary operations such as “plus”, “multiply”, “minus”, etc.
Feature Transformation Sequence: Assuming a dataset $\mathcal{D}=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{N}$ includes the original feature set $\mathcal{F}_{0}=\{f_{1},\ldots,f_{K}\}$ and predictive targets $y$. We transform the existing features using mathematical compositions $\tau$ consisting of feature ID tokens and operations to generate new and informative features. $K$ compositions are adopted to refine $\mathcal{F}_{0}$ to a better feature space $\tilde{\mathcal{F}_{0}}=\{\tilde{f}_{1},\ldots,\tilde{f_{K}}\}$. The collection of the $K$ compositions refers to the feature transformation sequence, which is denoted by $\Gamma=[\tau_{1},\cdots,\tau_{K}]$.
Postfix Expressions: The transformation sequence should be in a computable and machine-learnable format. The original infix representation has issues like redundancy, semantic sparsity, a high likelihood of illegal transformations, and an overly large search space. We introduce postfix expressions to solve these problems. Postfix expressions don’t need many brackets to determine calculation priority. Scanning from left to right suffices to reconstruct the corresponding sequence, greatly reducing sequence-modeling difficulty and computational cost. They also reduce the ambiguity of the transformation sequence. Most importantly, it reduces the search space from exponential to a finite set $|C|=|\mathcal{O}|+|\mathcal{F}_{0}|D+3$. Here, $|\mathcal{O}|$ represents the operation set size, $|\mathcal{F}_{0}|$ is the original feature set dimension, $D$ is feature numbers, and $3$ refers to start tokens <SOS><SOS>, separation token <SEP><SEP>, and end token <EOS><EOS>.
Figure 2. The framework of the Dataforge system.
3. Dataforge: Agentic Data Engineering
3.1. Overview of the Dataforge
Figure. 2 overviews Dataforge, which automates tabular data preparation and feature engineering via a structured, iterative pipeline. (1) Input and Cleaning: the input table is standardized through basic cleaning routines (schema/type normalization and missing-value handling) under preset modes (light, aggressive, or time-series). (2) Task-Level Routing: a lightweight router infers the learning setting (e.g., classification, regression, or unsupervised) and activates the corresponding feature-action space. (3) Action-Level Planning: conditioned on dataset profiles and past trajectories, an LLM planner proposes candidate feature operations (selection, transformation, and generation) from a tool library. (4) Grounding and Execution: each proposed action is validated by grounding checks (e.g., schema/type compatibility and logical constraints) before being executed by the backend. (5) Evaluation and Feedback: the evaluator measures task-specific metric (e.g., F1/AUC/RMSE) and drives two feedback loops: an inner loop rejects unsafe or invalid actions, while an outer loop performs iterative re-planning toward improved performance under a budget. (6) Experience and Reporting: the system records action trajectories, validation outcomes, and metric progression, enabling experience-guided planning and generating a human-readable report for traceability.
3.2. Hierarchical Routing
After data cleaning, Dataforge adopts a hierarchical routing strategy (Wang et al., 2025d; Bai et al., 2025a) to make decisions more accurate and efficient. Instead of directly searching over a large space of feature operations, we first perform a coarse-grained routing to a feature-operation family, and then conduct fine-grained planning within the selected family.
Task-level (coarse) routing: Here “task” refers to the data-engineering sub-task at the current iteration, i.e., choosing among three high-level operation families: feature selection, feature engineering, and feature transformation (Figure. 2). The router is rule-based and uses lightweight dataset profiles (e.g., feature dimensionality, type composition, missingness, cardinality, and simple distributional statistics) to decide which family is most promising to apply next. Intuitively, selection is preferred when the dataset exhibits strong redundancy or high dimensionality; transformation is preferred when scale/shape issues (e.g., heavy tails, outliers, or heterogeneous feature scales) dominate; and engineering is preferred when informative interactions or structured patterns are likely to exist. This coarse routing restricts subsequent planning to a smaller and more relevant candidate set, reducing confusion across operation types.
Action-level (fine) planning: Conditioned on the chosen family, an LLM-based planner selects concrete actions (e.g., specific selection criteria, transformation functions, or feature-generation methods) from the corresponding sub-library. By operating within a well-defined action subspace, the planner produces more consistent candidates and avoids irrelevant proposals, which improves routing accuracy and reduces unnecessary trials.
Why hierarchical routing: Let $\mathcal{A}$ denote the full tool library and $\mathcal{A}^{sel}$, $\mathcal{A}^{eng}$, $\mathcal{A}^{trans}$ denote the family-specific subsets. Hierarchical routing first chooses one subset and then searches within it, effectively shrinking the candidate space at each step. In our iterative setting, the outer optimization loop can revise the coarse choice in later iterations, enabling the system to alternate between families as needed.
3.3. Feature Engineering Tool Library
Dataforge organizes its feature operations into a family-structured tool library: feature selection, feature transformation, and feature engineering. Each tool corresponds to a parameterized feature action that can be proposed by the planner.
Feature selection: This family targets redundancy reduction and robustness by selecting informative subsets of columns. It includes standard filter and embedded operators (e.g., L1/LASSO, and tree-based importance), as well as group-wise selection primitives when features exhibit correlated blocks. Beyond classical operators, the library includes policy-based selectors inspired by reinforcement-learning-based selection (Liu et al., 2021a; Fan et al., 2021; Zhao et al., 2020; Wang et al., 2025e), which propose candidate subsets under budgeted evaluations.
Feature transformation: This family transforms existing features to improve numerical conditioning, handle scale and distribution mismatch, and mitigate noise/outliers, while preserving the original information content. We include common transformations such as normalization/standardization, power/log transforms, clipping for outliers, and discretization/binning. In addition to these hand-designed operators, the library can incorporate learned transformation modules that produce compact transformation programs or representations, leveraging reinforcement- or generation-based approaches to search over continuous or structured transformation spaces (Wang et al., 2023; Ying et al., 2024c; Gong et al., 2025d; Wang et al., 2022).
Feature engineering: This family constructs new features to expand representational capacity beyond the raw columns. We support widely used engineering primitives, including interaction/cross features (e.g., pairwise products and conjunctions), logical compositions, and group-wise aggregation patterns, which are effective when predictive signals emerge from multi-column interactions. For structured domains, the library can further incorporate domain-aware generators such as categorical hashing representations with hierarchical crossing and topology-aware feature space reconstruction for graph-structured attributes (Ying et al., 2025a, 2023), enabling systematic feature creation when hand-crafted templates are insufficient.
3.4. Dual Feedback Loops
We develop two collaborative feedback loops that turn the workflow into an adaptive, self-correcting process. The first loop ensures that every proposed action is executable (grounding for safety), while the second loop guides the system toward better downstream performance through iterative evaluation and re-planning.
Action Validation Loop for Safety: Given a candidate action proposed by the planner, Dataforge performs grounding checks before execution. These checks validate (i) schema compatibility (target columns exist and the action’s output schema is well-formed), (ii) type consistency (e.g., numeric operators are not applied to categorical strings; encoders receive valid category sets), and (iii) logical feasibility (e.g., avoiding division by zero, invalid logarithms, or degenerate transformations that collapse a column to constants). If any check fails, the action is rejected with an explicit failure reason, and the planner is prompted to revise the proposal. Only grounded actions proceed to execution, which significantly reduces runtime errors and prevents invalid pipelines from contaminating subsequent iterations.
Performance Optimization Loop for Adaptiveness: After a grounded action is executed, the evaluator measures downstream performance under the given task specification (e.g., F1 for classification and 1-RAE for regression). Dataforge maintains a best-so-far state that records the current best pipeline and its evaluation score. If the new action yields improvement beyond a small tolerance, the best state is updated and the action trajectory is stored as a positive experience; otherwise, the system rolls back to the best-so-far pipeline and triggers re-planning to explore alternative actions. This loop enables Dataforge to iteratively search over feature operations without human intervention, while remaining robust to noisy or non-monotonic improvements.
Stopping and budget control: Both loops are governed by user-defined budgets and early-stop criteria (e.g., maximum actions, minimum improvement, and time budget). The optimization terminates when the budget is exhausted or when no meaningful improvement is observed for several iterations, returning the best validated pipeline and its transformed dataset. Overall, the two loops form a closed-cycle mechanism that couples safe executability with evaluation-driven adaptivity.
3.5. Automatic Summary and Reporting
To ensure transparency and provenance, the system automatically logs the routing choices, all actions, grounding outcomes, machine learning performance metrics, and summarizes these logs into a structured, concise natural language report that outlines: (i) the cleaning and feature operations applied, (ii) the evolution of routing decisions, (iii) the validation checks performed, and (iv) the performance progression across iterations. The summary, viewable through the demo interface, translates internal reasoning into a human-readable narrative that lets non-experts trace data transformations and decisions, but also allows the feedback loop to collect experiences as benchmark data and perform self-reflection.
3.6. Pseudo-code of Dataforge
Algorithm 1 Dataforge (Compressed)
0:  Dataset $\mathcal{D}$; budget $\mathcal{B}$ (MaxActions, TimeBudget, MinImprove)
0:  Best transformed data $\mathcal{D}^{\star}$ and pipeline $\pi^{\star}$
1:  $\mathcal{D}^{\star}\leftarrow\textsc{Clean}(\mathcal{D})$; $\pi^{\star}\leftarrow[\,]$; $u^{\star}\leftarrow$-$\infty$
2:  for $t=1$ to MaxActions do
3:   if $\textsc{Stop}(\mathcal{B})$ then
4:    break
5:   end if
6:   $c\leftarrow\textsc{CoarseRoute}(\mathcal{D}^{\star})$ {Sel/Trans/Eng}
7:   $a\leftarrow\textsc{PlanAction}(c,\mathcal{D}^{\star})$;
8:   if $\neg\,\textsc{Ground}(a,\mathcal{D}^{\star})$ then
9:    continue
10:   end if
11:   $\mathcal{D}^{\prime}\leftarrow\textsc{Execute}(a,\mathcal{D}^{\star})$; $u\leftarrow\textsc{Evaluate}(\mathcal{D}^{\prime})$
12:   if $u>u^{\star}+\textsc{MinImprove}(\mathcal{B})$ then
13:    $\mathcal{D}^{\star}\leftarrow\mathcal{D}^{\prime}$; $\pi^{\star}\leftarrow\pi^{\star}\circ[a]$; $u^{\star}\leftarrow u$
14:   end if
15:  end for
16:  return $(\mathcal{D}^{\star},\pi^{\star})$
Algorithm 1 summarizes the overall workflow of Dataforge. Starting from a cleaned dataset, the agent iteratively (i) performs coarse routing to choose a feature-operation family, (ii) plans a concrete action, (iii) grounds the action to ensure executability, and (iv) executes and evaluates the updated data. The best-so-far pipeline is updated only when the downstream performance improves beyond a threshold, and the process terminates by budget or auto-stop.
Table 1. Downstream performance comparison on regression and classification datasets. For regression, we report 1-RAE (Relative Absolute Error), and for classification, F1-score (%).
Dataset German Credit Amazon Employee Ionosphere PimaIndian Messidor Feature SVMGuide3 OpenML 586 OpenML 618 Airfoil
Source UCIrvine Kaggle UCIrvine Kaggle UCIrvine LibSVM OpenML OpenML UCIrvine
Task Classification Classification Classification Classification Classification Classification Regression Regression Regression
Samples 1,000 32,769 351 768 1,151 1,243 1,000 1,000 1,503
Features 24 9 34 8 19 21 25 50 5
Original 74.20% 93.37% 93.37% 80.68% 69.09% 81.85% 0.6311 0.4402 0.5749
RDG 68.01% 92.31% 91.17% 76.04% 62.38% 78.68% 0.5681 0.3720 0.5193
LDA 63.91% 91.64% 65.53% 63.80% 47.52% 65.24% 0.1109 0.0521 0.2201
ERG 74.43% 92.43% 92.02% 76.17% 66.90% 82.62% 0.6147 0.3561 0.5193
NFS 68.67% 93.21% 91.17% 74.87% 63.77% 79.16% 0.5443 0.3473 0.5193
AFAT 68.32% 92.97% 92.87% 76.56% 66.55% 79.49% 0.5435 0.2472 0.5210
PCA 67.92% 92.29% 92.87% 63.80% 67.21% 67.60% 0.1109 0.1016 0.2730
TTG 64.51% 92.79% 90.31% 74.48% 66.46% 79.81% 0.5443 0.3467 0.5003
FeatLLM (Han et al., 2024) 76.35% 93.62% 95.38% 89.66% 72.62% 81.17% 0.6477 0.4597 0.5587
CAAFE (Hollmann et al., 2023) 59.92% 91.41% 92.84% 79.86% 66.10% 81.74% N/A N/A N/A
AutoFeat (Horn et al., 2019) 74.86% 93.29% 93.37% 80.86% 69.08% 82.54% 0.6329 0.4407 0.5746
OpenFE (Zhang et al., 2023) 74.50% 93.44% 93.37% 80.86% 69.09% 83.05% 0.6311 0.4402 0.5746
ELLM-FT (Gong et al., 2025c) 76.39% 93.17% 96.01% 89.66% 74.80% 82.70% 0.6328 0.4734 0.6174
Pure LLM (Llama 3.1) 75.43% 75.43% 96.68% 89.66% 75.61% 81.17% 0.7811 0.6653 0.6329
RL-Policy-1 (GRFG (Wang et al., 2022)) 68.29% 68.29% 93.16% 75.39% 69.24% 79.81% 0.5768 0.4562 0.5587
RL-Policy-2 (MOAT (Wang et al., 2023)) 72.44% 72.44% 95.69% 80.73% 73.02% 81.17% 0.6251 0.4734 0.5967
Dataforge 79.60% 94.41% 97.14% 90.14% 76.98% 84.64% 0.7849 0.7219 0.7492
4. Empirical Study
We demonstrate the functionality and effectiveness of Dataforge through both quantitative evaluation and visual inspection of its autonomous workflow. The goal is to show that Dataforge achieves reliable performance across diverse datasets while providing interpretable and user-friendly automation for non-expert users.
4.1. Setup
Datasets and protocol. We evaluate on diverse tabular benchmarks from UCIrvine (Public, 2023c), CPLM (Public, 2023a), Kaggle (Howard, 2023), and OpenML (Public, 2023b), covering both classification and regression (statistics in Table 1). All methods are evaluated with the same 5-fold cross-validation protocol, and we report the average score across folds.
Downstream evaluator and metrics. We assess feature engineering utility by training a fixed downstream model on the dataset with identical hyperparameters for all methods. For classification, we report macro-F1; for regression, we report $1$-RAE:
$1\text{-RAE}=1-\frac{\|\bm{y}_{pred}-\bm{y}_{real}\|_{1}}{\|\bm{y}_{real}-\bar{\bm{y}}_{real}\|_{1}}.$
Operator set. All applicable methods share the same operator set $\mathcal{O}=\mathcal{O}_{1}\cup\mathcal{O}_{2}$, where $\mathcal{O}_{1}=\{\texttt{sqrt},\texttt{square},\texttt{cube},\texttt{reciprocal},\texttt{log},\\ \texttt{sin},\texttt{cos},\texttt{tanh},\texttt{sigmoid},\texttt{standard},\texttt{normalize},\texttt{quantile}\}$ and $\mathcal{O}_{2}=\{+,-,\times,\div\}$.
LLM backbones. We use Llama-3.2-3B as our base LLM for the agent. We use a unified generation protocol with $\texttt{temperature}=0.7$, $\texttt{top\_p}=0.9$, and $\texttt{max\_new\_tokens}=500$.
Baselines and implementation. We compare against: (i) no/linear baselines, including Original, RDG, PCA (Maćkiewicz and Ratajczak, 1993), and LDA; (ii) classical automated feature engineering and search methods, including ERG, AFAT (Horn et al., 2020), AutoFeat (Horn et al., 2019), NFS (Chen et al., 2019), TTG (Khurana et al., 2018), and OpenFE (Zhang et al., 2023); (iii) RL-based policies, including GRFG (Wang et al., 2022) and MOAT (Wang et al., 2023); and (iv) LLM-based approaches, including FeatLLM (Han et al., 2024), CAAFE (Hollmann et al., 2023), ELLM-FT (Gong et al., 2025c), as well as a Pure LLM variant.
Configurations. All experiments were conducted on the Ubuntu 22.04.3 LTS operating system, with a 13th-generation Intel(R) Core(TM) i9-13900KF CPU and an NVIDIA GeForce RTX 4090 GPU. The experiments were conducted using Python 3.11.5 and PyTorch 2.0.1.
4.2. Study of Overall Performances.
Table 1 reports the downstream utility of different feature-engineering approaches on six classification and three regression benchmarks. We compare Dataforge with representative baselines from three families: classical automated feature engineering/search, RL-based policies, and LLM-based methods, as well as the Original features. The goal of this experiment is to evaluate whether an agentic, fully automated pipeline can consistently improve tabular representations across diverse datasets under a unified protocol.
Overall, Dataforge delivers the greatest and most consistent improvements across both classification and regression benchmarks. Compared with classical automated feature engineering and search methods (e.g., AutoFeat, AFAT, NFS, TTG, and OpenFE), Dataforge more reliably discovers beneficial transformations, suggesting that iterative, feedback-driven planning is better suited for navigating large operator spaces than one-shot heuristics or fixed search strategies. Compared with RL-based pipelines (GRFG/MOAT) and the Pure LLM variant, Dataforge exhibits substantially higher stability: it avoids low-utility or invalid proposals and maintains strong performance across heterogeneous datasets with different sizes and feature dimensions. These results indicate that coarse routing to an appropriate feature-operation family, followed by fine-grained action planning and grounding, effectively reduces candidate ambiguity and improves decision quality under a fixed budget.
Insights: First, the consistent improvements over Original features validate that Dataforge can autonomously refine tabular representations via a combination of selection, transformation, and feature construction, rather than relying on a single operator family. Second, the performance advantages are achieved without training a task-specific policy (unlike RL methods), making the approach readily deployable across datasets. Finally, Dataforge requires minimal user input (dataset and basic runtime constraints) while producing a fully automated pipeline, supporting the practicality of agentic data preparation as a general paradigm for tabular feature engineering.
4.3. Ablation Study
Table 2 ablates key components of Dataforge on three representative datasets (German Credit, Amazon Employee, and OpenML 618). For each variant, we report the downstream score (macro-F1 for classification and 1-RAE for regression), the valid-action rate Valid(%) (fraction of proposed actions that pass grounding and execute successfully), and the average number of evaluator calls, which reflects the interaction cost of the iterative pipeline.
Table 2. Ablation study of Dataforge on three representative datasets. Score is macro-F1 (%) for classification and 1-RAE for regression. Valid(%) reports the ratio of proposed actions that pass grounding and are executed successfully; Calls denotes the average number of evaluator calls.
German Credit (C) Amazon Employee (C) OpenML 618 (R)
Variant Score Valid(%) Calls Score Valid(%) Calls Score Valid(%) Calls
Full (Dataforge) 79.60% 98.51% 3 94.41% 99.90% 6 0.7219 98.84% 4
w/o Routing 76.86% 97.01% 2 93.96% 98.33% 3 0.7169 90.41% 3
w/o Grounding 77.49% 91.60% 5 94.36% 84.10% 11 0.7031 75.76% 5
w/o Perf Loop 76.27% 96.30% 1 93.95% 97.00% 1 0.6653 98.25% 1
[…]
Insights: These ablations highlight complementary roles of the components. Grounding primarily improves reliability by filtering non-executable actions, thereby reducing wasted attempts even when the final score is not dramatically affected. Hierarchical routing improves decision quality by narrowing the candidate space; without it, the agent is more likely to try low-utility actions and stop early under the same auto-stop rule. The performance optimization loop is crucial for refinement: iterative evaluation and best-so-far updates are needed to move beyond a first valid action and consistently reach the strongest pipeline under a fixed budget.
4.4. Study of Reliability and Automation.
We evaluate whether Dataforge can run end-to-end autonomously and remain reliable under iterative feature operations. In this study, we focus on two practical indicators: (i) execution reliability, measured by the rate of invalid attempts during action proposal and execution, and (ii) automation stability, reflected by the number of retries needed to obtain valid actions and complete a run.
(a) F1-score comparison.
(b) Error rate comparison.
Figure 3. Comparison of modular agent (DataAgent) and pure LLM baselines, showing higher reliability and fewer retries for the modular agent paradigm.
As shown in Figure 3, Dataforge exhibits consistently higher reliability than pure LLM baselines. Across all nine datasets, it completes the entire pipeline without interruption and typically converges with about two action retries on average. In contrast, directly using Llama 3.1 8B or Llama 3 8B to propose feature operations leads to a non-trivial failure rate (around 3–5%), indicating that purely generative pipelines are more likely to produce invalid or non-executable operations. Meanwhile, Dataforge achieves comparable or slightly better downstream utility (average score $\approx 0.783$), suggesting that improved reliability does not come at the expense of performance.
Insights: These results highlight the importance of the modular agent design. By combining hierarchical routing with grounding validation, Dataforge constrains the action space and filters out unsafe proposals before execution, thereby reducing retries and preventing workflow breakdowns. This makes the system more predictable for continuous operation and supports the goal of fully automated data preparation with minimal user intervention.
4.5. Study of Efficiency Trade-offs.
We further study the efficiency–performance trade-offs of Dataforge by comparing it with RL-based pipelines that explicitly optimize feature operations via training. We report three aspects of efficiency in Table 3: training cost (wall-clock training time), inference latency (time to complete a single run), and interaction cost measured by the number of calls (i.e., the number of evaluator/tool invocations needed to obtain a final result). This experiment aims to characterize whether an agentic pipeline can remain responsive while still achieving strong downstream utility.
Table 3. Efficiency comparison: time (seconds) and call times.
Metric RL-Policy-1 (GRFG (Wang et al., 2022)) RL-Policy-2 (MOAT (Wang et al., 2023)) Dataforge
Training Time (sec) 1493.7 649.4 0
Inference Time (sec) 0.5 2.14 3.89
Calls 13 22 2
Performance 0.5768 0.6251 0.7832
Table 3 shows that Dataforge eliminates the dominant overhead of RL-style optimization. Both RL-Policy-1 (GRFG) and RL-Policy-2 (MOAT) require substantial training time (1493.7 s and 649.4 s), whereas Dataforge shows zero training cost and can be directly applied to new datasets. At inference time, Dataforge completes a full workflow in 3.89 s on average, which is slightly slower than RL inference (0.5–2.14 s) but remains lightweight in absolute terms. Notably, it achieves this with only two calls on average, compared to 13–22 calls for the RL baselines, indicating a markedly lower interaction/sample complexity under the same evaluation setting. Despite the reduced call budget, Dataforge attains the best downstream performance (0.7832), suggesting that its iterative planning is highly cost-effective.
Insights: These results highlight a favorable operating point for practical deployment. While RL policies can achieve fast inference after training, the up-front training time and higher call counts make them less attractive when datasets change frequently or when rapid, on-demand data preparation is required. In contrast, Dataforge trades a small increase in per-run latency for eliminating training altogether and substantially reducing interaction cost, yielding strong performance with high responsiveness. This supports the use of modular agentic feature engineering as a practical alternative to training-heavy optimization methods.
5. Demo and Case Studies
5.1. User Interface
Figure 4. The interface of Dataforge.
Figure. 4 illustrates the user interface of Dataforge, which offers an intuitive environment for autonomous tabular data transformation. Users can upload datasets in standard formats (.csv, .parquet, and .hdf), label predictors and response column in a machine learning task, and configure runtime parameters, such as, Max Actions, Min Improvement, and Time Budget. After uploading datasets, users can select from pre-loaded examples. During execution, the interface visualizes real-time logs of thinking steps in reasoning, data cleaning and feature engineering actions, and predictive performance metrics (e.g., F1-score, AUC, MAE, and RMSE depending on classification, regression, and unsupervised tasks ), so that users can observe how the data agent progressively refines data quality. Upon completion, it summarizes all executed actions, validation outcomes, and final performance statistics, to provide transparency and provenance, as well as collect action historical trajectories for benchmark data of training data agents. The interface is designed to be modular and extensible, supporting future integration with interactive dashboards and collaborative agent environments.
5.2. Toward Accurate and Efficient Heart-Disease Detection
Accurate identification of heart disease is one of the central challenges in clinical cardiology. Modern diagnostic models rely on dozens of imaging and physiological indicators, but these measurements are often noisy, redundant, or inconsistent across patients. As a result, even experienced cardiologists face difficulty in building reliable predictive models from raw data, and existing computational tools usually require manual data engineering and expert parameter tuning.
Figure 5. Dataforge Dealing with Heart-Disease Detection.
To address this challenge, DataForge is applied to an automated diagnostic classification task using the SPECTF Heart dataset, which contains patient-level imaging signals used for detecting abnormal cardiac function. The system autonomously analyzes the uploaded dataset, identifies the most informative diagnostic indicators, and removes irrelevant or conflicting variables. It then constructs refined combinations of cardiac features that better represent patient conditions, iteratively evaluating improvements until no further diagnostic gain can be achieved.
Throughout this process (Figure. 5), no manual intervention or algorithmic knowledge is needed. The workflow completes in about 100 seconds and generates a concise report and a new dataset that achieves higher diagnostic accuracy with fewer features, reducing the total number of variables from 44 to 20 while improving the predictive score from 0.772 to 0.840. This demonstrates that DataForge can serve as an intelligent assistant for cardiology, automatically transforming complex measurements into a cleaner, smaller, and more reliable dataset for accurate and efficient heart-disease detection.
5.3. Chemical Design
Figure 6. Dataforge for materials discovery.
Figure 6 illustrates how Dataforge supports a practical materials discovery workflow. Starting from a small seed set of physical experiments, we obtain an initial tabular dataset with multiple target properties. Dataforge then enhances the dataset via automated feature operations and produces an AI-ready version together with traceable logs and a natural-language report. Training a downstream property predictor on the enhanced data improves predictive utility across targets, which enables more reliable screening and prioritization of candidate formulations, thereby reducing the number of costly follow-up experiments.
6. Related Work
6.1. Automated Feature Engineering
Feature transformation. Prior work formulates feature transformation as searching over composition trajectories of operators and operands. Early systems suggest generating a large candidate pool followed by selection to control redundancy (Katz et al., 2016). To improve trajectory quality, many methods adopt structured discrete search, including heuristic-guided exploration and other discrete optimization strategies (Kanter and Veeramachaneni, 2015; Khurana et al., 2016; Tran et al., 2016), as well as evolutionary algorithms for population-based exploration (Zhu et al., 2022a; Gong et al., 2024). Reinforcement learning further models transformations as sequential decision making and learns policies from downstream rewards (Wang et al., 2022; Ying et al., 2025b; Bai et al., 2025b), with traceable/evolutionary variants emphasizing controllable search dynamics and interpretability (Xiao et al., 2023a, 2024). Beyond discrete search, continuous optimization methods embed feature programs into differentiable latent spaces and apply gradient-based updates (Wang et al., 2023; Zhu et al., 2022b; Gong et al., 2025b), while facing the challenge of mapping optimized embeddings back to valid executable transformations.
Feature selection. Classical feature selection includes similarity/spectral criteria (e.g., Laplacian Score (He et al., 2005), SPEC (Zhao and Liu, 2007), Fisher Score (Duda et al., 2006), Trace Ratio (Nie et al., 2010)), information-theoretic filters (e.g., Information Gain (Lewis, 1992), mutual information (Battiti, 1994), mRMR (Peng et al., 2005), FCBF (Yu and Liu, 2003)), sparse learning (e.g., $\ell_{p}$-regularized methods (Tibshirani, 1996; Zhu et al., 2003)), and statistics-based measures (e.g., T-test (Davis et al., 1986), Chi-square (Liu and Setiono, 1995), Gini (Gini, 1912), CFS (Hall and Smith, 1999)). Recent work leverages deep and RL formulations to improve scalability and search quality, including deep feature selection (Xiao et al., 2023b; Ying et al., 2024a, b), RL-based selection (Wang et al., 2025e), Monte-Carlo style methods (MARFS (Liu et al., 2023), MCRFS (Liu et al., 2021b)), and specialized settings such as incomplete multi-view unsupervised selection (Huang et al., 2023).
LLMs for feature engineering. Recent studies explore using LLMs to propose feature transformations or engineered features via natural-language reasoning, enabling broader operator coverage and reduced manual effort (Han et al., 2024; Hollmann et al., 2023; Gong et al., 2025c, d, a; Wang et al., 2025c; Xie et al., 2025). However, purely generative pipelines can suggest invalid or non-executable actions and often require additional validation and iterative control to ensure reliability and stable improvements.
6.2. Agentic Platform
Agentic systems augment LLMs with multi-step reasoning, planning, and tool use (Zheng et al., 2026; Wang et al., 2025a; Xie and Ju, 2026), enabling them to iteratively act in an environment rather than producing one-shot outputs (Yin et al., 2025). A representative paradigm is ReAct, which interleaves reasoning traces with action/tool calls to improve task completion and error recovery (Yao et al., 2022).
Building on general tool-using agents, web agents study how LLMs perceive and interact with complex web environments (e.g., multi-page, multi-tab interfaces) to accomplish real-world browsing tasks. Benchmarks such as WebArena provide realistic web interaction environments for evaluating such agents (Zhou et al., 2023), and systems like WebVoyager demonstrate end-to-end web navigation and action execution at scale (He et al., 2024).
Recently, the same agentic principles have been extended to the data domain. Fu et al. propose autonomous data agents that combine task decomposition, action reasoning, grounding into executable code/tool calls, and adaptive workflow planning to automate data-to-knowledge pipelines (Fu et al., 2025). This perspective highlights new opportunities and challenges for trustworthy, scalable data operations (e.g., preprocessing, selection, transformation, and repairs) and motivates agentic platforms like ours that emphasize grounded execution and iterative feedback for reliable data engineering (Fu et al., 2025).
7. Conclusion
We presented Dataforge, an LLM-powered agentic data engineering platform for tabular data. It is automatic, safe, and non-expert friendly. By combining hierarchical routing with evaluation-driven refinement and automatic stopping, Dataforge transforms raw datasets into AI-ready representations end-to-end. Experiments on nine public benchmarks show strong and robust gains over classical AutoFE/search, RL-based, and LLM baselines, while ablations highlight the roles of routing, grounding, and iterative optimization.